OpenStudy (anonymous):

Evaluate this expression. 6/6 \ 2/3

OpenStudy (anonymous):

will you please elaborate ?

OpenStudy (anonymous):

6/6 divided by 2/3

OpenStudy (anonymous):

when dividing fraction by another fraction just copy the first factor and get the reciprocal of the second fraction AND change the division process to multiplication process :) 6/6 * 3 /2 now can you have it by yourself ?

OpenStudy (anonymous):

6*3 6*2 18/12= 9/6=3/2 3/2??